# Noise-Robust Hearing Aid Voice Control

### v2

@@ II-A Hearing Aid Transfer Function Measurement Setup
Own-voice transfer functions (OVTFs) and head-related transfer functions (HRTFs) were estimated from the experimental setup depicted in Fig. 1. One at a time, a total of 29 subjects sat in a soundproof sound studio while wearing HA prototypes [^3] For experimental purposes, and assuming symmetry, we just consider the left ear HA, as shown in Fig. 1. by Demant A/S that embed 2 (front and rear) BTE microphones and 1 IEC microphone. Subjects were surrounded by 16 equidistantly-spaced loudspeakers laid out in a circular array at eye-height.
@@ II-B Noisy Own-voice Generation Model
where $\alpha$ is a noise scaling factor.
@@ II-C Dataset Creation and Structure
Mouth reference microphone signals $x(n)$ were obtained from the GSCD v2 [8], which is composed of more than 100k 1-second long utterances comprising one word each. Similar to [17], we first extracted clean training, validation and test sets out of the GSCD by picking all the signals in each set with an a posteriori SNR above 40 dB [^4] Given a GSCD utterance, to estimate its a posteriori SNR, its first 100 ms were used to calculate the noise power, whereas the rest was employed to compute the noisy speech power. The a posteriori SNR is defined as the ratio of the two.. Twelve classes (which are approximately balanced across the training, validation and test sets) were defined: the 10 standard keywords of the GSCD (“yes”, “no”, “up”, “down”, “left”, “right”, “on”, “off”, “stop” and “go”), the filler class (comprised of the other 25 GSCD words), and ambient noise (i.e., no own voice is present).
While we were provided with transfer functions from 22 out of the 29 subjects who took part in the measurements (see Subsec. II-A), only 5 of the subjects showed good ear occlusion and were kept for noisy own-voice generation. From those 5, 3 were allocated to the training set, 1 more, to the validation set, and the final one, to the test set. In turn, each GSCD training speaker was randomly assigned 1 out of the 3 training HA subjects.
@@ IV Experiments and Results
Finally, note from Fig. 4 that KWS performance increases with the number of model parameters and starts to stagnate from 50k parameters. This makes BC-ResNet-3 (which has 54,168, 55,368 and 56,568 parameters for 1-, 2- and 3-microphone inputs, respectively) a good candidate to be deployed in HAs taking into account that they are low-resource devices. Table II shows detailed KWS accuracy results (broken down by SNR and seen/unseen noises) and real-time factors (RTFs) [^5] RTFs, defined as the ratio between the time taken to process an input and the input duration [24], were measured on an Intel Xeon Silver 4214R CPU with a clock frequency of 2.4 GHz. obtained by BC-ResNet-3 employing the previously tested microphones and combinations of them. From this table, we can clearly assess how the combination of noise-robust BCS with the broader-bandwidth BTE microphone signals yields a strong voice control performance at extremely challenging SNRs.
@@ V Conclusion
In this work, we have aimed at establishing a baseline for noise-robust HA voice control. This has been accomplished through 1) the creation of a publicly available11footnotemark: 1 HA noisy dataset comprising both regular air- and bone-conducted own voice, and 2) the evaluation of light, state-of-the-art KWS models for HA voice control using such a dataset. Experimental results have demonstrated that it is possible to achieve robust voice control by exploiting BCS from an IEC microphone as well as voice control performance can be boosted by adding complementary noisy higher frequency information from BTE microphones.

### v3

@@ II-A Hearing Aid Transfer Function Measurement Setup
Own-voice transfer functions (OVTFs) and head-related transfer functions (HRTFs) were estimated from the experimental setup depicted in Fig. 1. One at a time, a total of 22 subjects sat in a soundproof sound studio while wearing HA prototypes [^3] For experimental purposes, and assuming symmetry, we just consider the left ear HA, as shown in Fig. 1. by Demant A/S that embed 2 (front and rear) BTE microphones and 1 IEC microphone. Subjects were surrounded by 16 equidistantly-spaced loudspeakers laid out in a circular array at eye-height.
@@ II-B Noisy Own-voice Generation Model
where $\alpha$ is a noise scaling factor. Note that we do not account for the Lombard effect despite it is known to have an impact on speech recognition performance [16].
@@ II-C Dataset Creation and Structure
Mouth reference microphone signals $x(n)$ were obtained from the GSCD v2 [8], which is composed of more than 100k 1-second long utterances comprising one word each. Similar to [18], we first extracted clean training, validation and test sets out of the GSCD by picking all the signals in each set with an a posteriori SNR above 40 dB. Twelve classes (which are approximately balanced across the training, validation and test sets) were defined: the 10 standard keywords of the GSCD (“yes”, “no”, “up”, “down”, “left”, “right”, “on”, “off”, “stop” and “go”), the filler class (comprised of the other 25 GSCD words), and ambient noise (i.e., no own voice is present).
While we were provided with transfer functions from 22 subjects (see Subsec. II-A), only 5 of the subjects showed good ear occlusion (which was determined by visual inspection of HRTFs) and were kept for noisy own-voice generation. From those 5, 3 were allocated to the training set, 1 more, to the validation set, and the final one, to the test set. In turn, each GSCD training speaker was randomly assigned 1 out of the 3 training HA subjects.
@@
Roughly, the training, validation and test sets contain 18.8 hrs, 2.1 hrs and 2.4 hrs, respectively, of multi-channel audio.
@@ IV Experiments and Results
Finally, note from Fig. 4 that KWS performance increases with the number of model parameters and starts to stagnate from 50k parameters. This makes BC-ResNet-3 (which has 54,168, 55,368 and 56,568 parameters for 1-, 2- and 3-microphone inputs, respectively) a good candidate to be deployed in HAs taking into account that they are low-resource devices. Table II shows detailed KWS accuracy results (broken down by SNR and seen/unseen noises) and real-time factors (RTFs) [^4] RTFs, defined as the ratio between the time taken to process an input and the input duration [25], were measured on an Intel Xeon Silver 4214R CPU with a clock frequency of 2.4 GHz. obtained by BC-ResNet-3 employing the previously tested microphones and combinations of them. From this table, we can clearly assess how the combination of noise-robust BCS with the broader-bandwidth BTE microphone signals yields good voice control performance at extremely challenging SNRs.
@@ V Conclusion
In this work, we have aimed at establishing a baseline for noise-robust HA voice control. This has been accomplished through 1) the creation of a publicly available22footnotemark: 2 HA noisy dataset comprising both regular air- and bone-conducted own voice, and 2) the evaluation of light, state-of-the-art KWS models for HA voice control using such a dataset. Experimental results have demonstrated that it is possible to achieve robust voice control by exploiting BCS from an IEC microphone as well as voice control performance can be boosted by adding complementary noisy higher frequency information from BTE microphones.